Joe has a collection of nickels and dimes that is worth $9.60. If the number of nickels was tripled and the number of dimes was decreased by 31, the value of the coins would be $6.70. How many nickels and dimes does he have?
Answer by mananth(16946)

Let the nickels be x
Dimes = y
5x+10y=960---------------(1)
5(3x)+10(y-31)=670
15x+10y-310=670
15x+10y=980 ---------------(2)
Multiply equation (1) by -3
we get
-15x-30y=-2880
Add this equation to (2)
-20y=-1900
/-20
y=95 ---- dimes
plug x in any equation you get x=2 nickels
CHECK
5*2+10*95=960